# FEALPy v3: A Cross-platform Intelligent Numerical Simulation Engine

### v1

@@ 4 Numerical Examples
In discretization, the domain $\Omega$ was initially partitioned into a uniform $4\times 4\times 4$ mesh of tetrahedrons, and was successively uniformly refined four times. Finite element spaces $V_{h}\subset H^{1}(\Omega)$ was constructed using linear or quadratic Lagrange elements on each mesh level. Lst. 6 shows the construction of the finite element space.
@@ 3 LagrangeFESpace, TensorFunctionSpace
Listing 6: A tensor function space with DoFs shaped (2,) was constructed on a tetrahedral mesh.
@@
where $V_{h}^{0}$ is the test space satisfying homogeneous Dirichlet boundary conditions. See Lst. 7 for the code. The non-homogeneous Dirichlet conditions are enforced in a strong form.
@@ 11 LinearElasticIntegrator(material=material)
Listing 7: Construct a bilinear form based on the space, and integrators can be added to the form to define the weak form. The assembly method returns a sparse matrix. The construction of linear forms is similar.
@@
The computed $L^{2}$ errors and the corresponding mesh sizes are summarized in Tab. 5.
h 1/4 1/8 1/16 1/32 1/64
$p=1$ DoF 375 2187 14739 107811 823875
$\|{\bm{u}}_{\text{exact}}-{\bm{u}}_{h}\|_{L^{2}(\Omega)}$ 3.0271e-2 1.1616e-2 3.5874e-3 9.8912e-4 2.5827e-4
Order - 1.3818 1.6951 1.8587 1.9372
$p=2$ DoF 2187 14739 107811 823875 -
$\|{\bm{u}}_{\text{exact}}-{\bm{u}}_{h}\|_{L^{2}(\Omega)}$ 5.0775e-3 6.0218e-4 6.8054e-5 7.9984e-6 -
Order - 3.0759 3.1454 3.0889 -
Table 5: Convergence results for the linear elasticity problem. The error decreases systematically as the mesh is refined.
The calculated convergence rate approaches the optimal theoretical rate 2 for linear elements, and 3 for quadratic elements in the $L^{2}$-norm, which strongly validates the correct implementation of the FEM for linear elasticity within FEALPy and confirms its numerical precision.
@@
We use the $C^{1}$-conforming finite element space proposed by [CHENLONG2021, HUJUN2023], with shape functions of degree $k\geq 9$ on tetrahedral meshes.
One of FEALPy’s key features is its modular design, which allows users to implement new finite elements by defining only the local basis functions and degrees of freedom, while reusing existing infrastructure for mesh handling, assembly, and linear algebra. To solve this problem, we simply replaced the standard Lagrange element with the smooth element ($C^{m}$ Conforming space, implemented by [CHEN2025]), on a initially $1\times 1\times 1$ uniform mesh of tetrahedrons. See Lst. 8.
@@
Listing 8: The Smooth Finite Element can be implemented by introducing a new function space and integrator.
Tab. 6 shows the convergence on a sequence of uniformly refined tetrahedral meshes, which confirm optimal convergence rates in the $L^{2}$-norm, as theoretically expected.
@@ 4.3 Burgers’ Equation with Moving Mesh
In FEALPy, we have implemented a high-dimensional, vectorized moving mesh algorithm based on the Harmonic map[RUOLI2001, RUOLI2002] and the MMPDE gradient flow approach[HUANG2001]. A MMesher class is provided to control the moving mesh process, as presented in Lst. 9.
@@
Listing 9: MMesher is designed to control the moving mesh process. It is irrelevant to mesh types and FE spaces.
@@
We apply the forward Euler time scheme with $a=0.005$ and time step $\Delta t=0.001$. The initial moving meshes include uniform triangular grids of sizes $30\times 30$, $45\times 45$, and $60\times 60$, as well as a $45\times 45$ quadrilateral mesh. For comparison, a fixed uniform triangular grid of size $160\times 160$ is used. See Tab. 7 for $L^{2}$ errors and times, Fig. 6 and 7 for the solution and moving mesh respectively.
Type and Density $t=0.5$ $t=0.75$ $t=1.0$ $t=1.25$ $t=1.5$ Wall Time (s)
@@
In the second test, we compare computational performance between CPU and GPU execution. With $a=0.005$ and $\Delta t=0.001$ over the time interval $[0,0.25]$, we measure the total runtime on triangular meshes of sizes $60\times 60$, $80\times 80$, $120\times 120$, and $160\times 160$. The results, summarized in Tab. 8, show that the PyTorch-GPU backend achieves the best performance.
@@ 4.4 Inverse Problem with Learning Automated FEM
Inspired by DDSM[GUO2021], a novel operator learning framework termed $\gamma$-deepDSM[ZHENG2024] was proposed integrating a learnable preconditioning operator with a neural network for enhanced reconstruction. The method consists of two synergistic components (as shown in Fig. 8):
@@
    Preprocessing via Learnable Fractional Laplace-Beltrami Operator [MATEUSZ2017]. Boundary data $\xi_{l}=g_{D,l}-\Lambda_{\sigma_{0}}g_{N,l}$ is transformed using an fLB operator with learnable fractional orders $\gamma_{l}$:
@@
    are handled as a tensor and fed into a U-Net[RONNEBERGER2015] to reconstruct the conductivity image $\sigma$.
Thanks to the PyTorch support of FEALPy, we can easily implement the $\gamma$-deepDSM framework by directly embedding FEM code into NN modules, as shown in Lst. 10.
@@
Listing 10: A PyTorch NN module containing a FEM solver. The ’self.solver’ here is a wrapper for FEALPy code.
@@
This example highlights FEALPy’s unified tensor-based architecture, which seamlessly integrates traditional FEM with DL workflows. The work creates a powerful framework for solving challenging inverse problems while maintaining mathematical rigor and computational efficiency.
@@ 4.5 Path Planning with Metaheuristic Optimization
A grid-based AGV (Automated Guided Vehicle) path planning problem is considered on a $20\times 20$ map, where cells with value 1 represent obstacles and 0 represent free space. Given the start point $(0,0)$ and end point $(19,19)$, we solve the problem using the Particle Swarm Optimization (PSO)[HAN2023] algorithm implemented in FEALPy (Lst. 11). The resulting path is shown in Fig. 10.
@@ 3 PathPlanningModelManager,
Listing 11: 2D path planning model using PSO in FEALPy.OPT. Note that the PathPlanningModelManager generates conputational models (such as the planner in this code) as wrappers for the path planning scripts, which improved the readability and usability of the code in basic teaching. Readers can examine its internal implementation to learn how to use the ParticleSwarmOpt algorithm directly.
@@
A 3D terrain environment is constructed from the dataset ChrismasTerrain.tif, with several cylindrical threat regions to be avoided. Given the start point $(800,100,150)$ and end point $(100,800,150)$, the path was optimized using the Snow Ablation Optimization (SAO)[DENG2023] algorithm implemented in FEALPy. Code for this problem is similar, but with a different map type (the thread regions and terrain data), as shown in Lst. 12. The optimized trajectory is shown in Figure  10, which shows the 3D terrain surface and feasible paths.
@@
Listing 12: 3D terrain-based path planning using SAO in FEALPy.OPT. The parts that are the same as in the previous example have been omitted.
@@ 5.1 SOPTX: High-Performance Multi-Backend Topology Optimization
The SOPTX[HE2025] framework was developed based on FEALPy primarily targeting applications in structural topology optimization. SOPTX fully inherits and extends FEALPy’s multi-backends feature, diverse numerical algorithm components, and general-purpose handling capabilities for mesh and geometry. Key technologies such as GPU parallel acceleration, sensitivity analysis based on AD, and high-performance matrix assembly techniques are further integrated based on this foundation, which significantly enhances the computational efficiency of topology optimization problems while ensuring flexibility and extensibility.
@@ 5.2 FractureX: Flexible and Efficient Phase-Field Fracture Simulation
The core strength of FractureX lies in its modular architecture, which cleanly decouples key physical and numerical components. This design allows users to effortlessly combine different strain energy decomposition models (e.g., isotropic, anisotropic, spectral, and hybrid models [MIEHE2010]), energy degradation functions, and crack surface density functions (AT1, AT2 [WU2018]). Such modularity facilitates rapid prototyping of novel constitutive models and direct comparison between different modeling approaches within a unified codebase.
FractureX adopts an automated adaptive mesh refinement (AMR) strategy driven by a recovery-type error estimator[TIAN2026], which significantly reduce the cost and improve the accuracy through dynamically focusing the mesh resolution on the fracture region. The main solver includes a robust Newton-Raphson scheme for the strongly coupled displacement-phase-field system. Inherited from FEALPy, GPU acceleration is ready for large-scale 3D simulations, and AD is also available to skip conventional matrix assembly in rapid tests for new models.
@@
6 Conclusions
This work introduced FEALPy, a tensor-centric numerical simulation engine that unifies traditional numerical simulation methods with modern tensor computation/DL frameworks. By organizing computations around a consistent tensor abstraction layer, FEALPy achieves interoperability across multiple backends such as NumPy, PyTorch, and JAX, allowing seamless execution on diverse hardware architectures including CPUs and GPUs.
The modular architecture of FEALPy—comprising the tensor, commonality, algorithmic, and field layers—effectively separates mathematical abstraction from computational realization. This design provides both clarity and flexibility for researchers to implement, extend, and combine numerical algorithms without modifying the underlying data structures. The framework further supports AD through nearly all computational stages, establishing a direct interface between classical PDE solvers and differentiable programming paradigms.
@@ 6 Conclusions
Source code of FEALPy is available at https://github.com/weihuayi/fealpy. And the source code of numerical examples are available at https://github.com/weihuayi/fealpy/tree/master/example. The user manual for FEALPy: http://suanhai.com.cn/intro.html.

### v2

@@ 4 Numerical Examples
In discretization, the domain $\Omega$ was initially partitioned into a uniform $4\times 4\times 4$ mesh of tetrahedrons, and was successively uniformly refined four times. Finite element spaces $V_{h}\subset H^{1}(\Omega)$ was constructed using linear or quadratic Lagrange elements on each mesh level. Listing 7 shows the construction of the finite element space.
@@ 3 LagrangeFESpace, TensorFunctionSpace
Listing 7: A tensor function space was constructed on a tetrahedral mesh. The ’shape’ parameter specifies the tensor shape and DoF arrangement.
@@
where $V_{h}^{0}$ is the test space satisfying homogeneous Dirichlet boundary conditions. See Listing 8 for the code. The non-homogeneous Dirichlet conditions are enforced in a strong form.
@@ 11 LinearElasticIntegrator(material=material)
Listing 8: Construct a bilinear form based on the space, and integrators can be added to the form to define the weak form. The assembly method returns a sparse matrix. The construction of linear forms is similar.
@@
The computed $L^{2}$ errors, order of convergence and the corresponding mesh sizes are summarized in Table 5 and Table 6, where the linear systems were solved by Pardiso.
p h 1/2 1/4 1/8 1/16 1/32 1/64
1 DoF - 375 2187 14739 107811 823875
$L^{2}$ Error - 3.0271e-2 1.1616e-2 3.5874e-3 9.8912e-4 2.5827e-4
Order - - 1.3818 1.6951 1.8587 1.9372
2 DoF - 2187 14739 107811 823875
$L^{2}$ Error - 5.0775e-3 6.0218e-4 6.8054e-5 7.9984e-6
Order - - 3.0759 3.1454 3.0889
4 DoF 2187 14739 107811 823875
$L^{2}$ Error 1.3435e-03 6.0182e-05 2.1914e-06 7.2287e-08
Order - 4.4805 4.7794 4.9219
8 DoF 14739 107811 823875
$L^{2}$ Error 2.2096e-06 4.9601e-09 9.9604e-12
Order - 8.7992 8.9600
Table 5: Convergence results for the linear elasticity problem. The $L^{2}$-norm error decreases systematically as the mesh is refined.
p h 1 1/2 1/4
10 DoF 3993 27783 206763
$L^{2}$ Error 1.5787e-14 6.5569e-14 9.3363e-14
Order - -2.0543 -0.5098
Table 6: Convergence results for the linear elasticity problem with $p=10$. This is an exact solution up to machine precision and the error of linear system solver.
The calculated convergence order approaches the optimal theoretical order, (e.g. 2 for linear elements, and 3 for quadratic elements in the $L^{2}$-norm) strongly validating the correct implementation of the FEM for linear elasticity within FEALPy and confirms its numerical precision.
@@ 10 DoF 3993 27783 206763
We use the $C^{1}$-conforming finite element space proposed by [10, 17], with shape functions of degree $k\geq 9$ on tetrahedral meshes.
Benefiting from FEALPy’s modular design, users are allowed to implement new finite elements by defining only the local basis functions and degrees of freedom, while reusing existing infrastructure for mesh handling, assembly, and linear algebra. To solve this problem, we simply replaced the standard Lagrange element with the smooth element ($C^{m}$ Conforming space, implemented by [8]), on a initially $1\times 1\times 1$ uniform mesh of tetrahedrons. See Listing 9.
@@
Listing 9: The Smooth Finite Element can be implemented by introducing a new function space and integrator.
Table 7 shows the convergence on a sequence of uniformly refined tetrahedral meshes, confirming optimal convergence rates in the $L^{2}$-norm, as theoretically expected.
@@ 4.3 Burgers’ Equation with Moving Mesh
In FEALPy, we have implemented a high-dimensional, vectorized moving mesh algorithm based on the Harmonic map[24, 25] and the MMPDE gradient flow approach[19]. A MMesher class is provided to control the moving mesh process, as presented in Listing 10.
@@
Listing 10: MMesher is designed to control the moving mesh process. It is irrelevant to mesh types and FE spaces.
@@
We apply the forward Euler time scheme with $a=0.005$ and time step $\Delta t=0.001$. The initial moving meshes include uniform triangular grids of sizes $30\times 30$, $45\times 45$, and $60\times 60$, as well as a $45\times 45$ quadrilateral mesh. For comparison, a fixed uniform triangular grid of size $160\times 160$ is used. See Table 8 for $L^{2}$ errors and times, Figure 6 and 7 for the solution and moving mesh respectively.
Type and Density $t=0.5$ $t=0.75$ $t=1.0$ $t=1.25$ $t=1.5$ Time (s)
@@
In the second test, we compare computational performance between CPU and GPU execution. With $a=0.005$ and $\Delta t=0.001$ over the time interval $[0,0.25]$, we measure the total runtime on triangular meshes of sizes $60\times 60$, $80\times 80$, $120\times 120$, and $160\times 160$. The results, summarized in Table 9, show that the PyTorch-GPU backend achieves the best performance.
@@ 4.4 Inverse Problem with Learning Automated FEM
Inspired by DDSM[13], a novel operator learning framework termed $\gamma$-deepDSM[37] was proposed integrating a learnable preconditioning operator with a neural network for enhanced reconstruction. The method consists of two synergistic components (Figure  8):
@@
    Preprocessing via Learnable Fractional Laplace-Beltrami Operator [23]. Boundary data $\xi_{l}=g_{D,l}-\Lambda_{\sigma_{0}}g_{N,l}$ is transformed using an fLB operator with learnable fractional orders $\gamma_{l}$:
@@
    are handled as a tensor and fed into a U-Net[31] to reconstruct the conductivity image $\sigma$.
Thanks to the PyTorch support of FEALPy, we can easily implement the $\gamma$-deepDSM framework by directly embedding FEM code into NN modules, as shown in Listing 11.
@@
Listing 11: A PyTorch NN module containing a FEM solver. The ’self.solver’ here is a wrapper for FEALPy code.
@@
This example highlights FEALPy’s unified tensor-based architecture that seamlessly integrates traditional FEM with DL workflows. The work creates a powerful framework for solving challenging inverse problems while maintaining mathematical rigor and computational efficiency.
@@ 4.5 Path Planning with Metaheuristic Optimization
A grid-based AGV (Automated Guided Vehicle) path planning problem is considered on a $20\times 20$ map, where cells with value 1 represent obstacles and 0 represent free space. Given the start point $(0,0)$ and end point $(19,19)$, we solve the problem using the Particle Swarm Optimization (PSO)[14] algorithm implemented in FEALPy (Listing 12). The resulting path is shown in Figure 10.
@@ 3 PathPlanningModelManager,
Listing 12: 2D path planning model using PSO in FEALPy.OPT. Note that the PathPlanningModelManager generates conputational models (such as the planner in this code) as wrappers for the path planning scripts, which improved the readability and usability of the code in basic teaching. Readers can examine its internal implementation to learn how to use the ParticleSwarmOpt algorithm directly.
@@
A 3D terrain environment is constructed from the dataset ChrismasTerrain.tif, with several cylindrical threat regions to be avoided. Given the start point $(800,100,150)$ and end point $(100,800,150)$, the path was optimized using the Snow Ablation Optimization (SAO)[12] algorithm implemented in FEALPy. Code for this problem is similar, but with a different map type (the thread regions and terrain data), as shown in Listing 13. The optimized trajectory in Figure 10 shows the 3D terrain surface and feasible paths.
@@
Listing 13: 3D terrain-based path planning using SAO in FEALPy.OPT. The parts that are the same as in the previous example have been omitted.
@@ 5.1 SOPTX: High-Performance Multi-Backend Topology Optimization
The SOPTX[16] framework was developed based on FEALPy primarily targeting applications in structural topology optimization. SOPTX fully inherits and extends FEALPy’s multi-backends feature, diverse numerical algorithm components, and general-purpose handling capabilities for mesh and geometry. Key technologies such as GPU parallel acceleration, sensitivity analysis based on AD, and high-performance matrix assembly techniques are further integrated based on this foundation, which significantly enhances the computational efficiency of topology optimization problems while ensuring flexibility and extensibility.
@@ 5.2 FractureX: Flexible and Efficient Phase-Field Fracture Simulation
The core strength of FractureX lies in its modular architecture, which cleanly decouples key physical and numerical components. This design allows users to effortlessly combine different strain energy decomposition models (e.g., isotropic, anisotropic, spectral, and hybrid models [26]), energy degradation functions, and crack surface density functions (AT1, AT2 [36]). Such modularity facilitates rapid prototyping of novel constitutive models and direct comparison between different modeling approaches within a unified codebase.
FractureX adopts an automated adaptive mesh refinement (AMR) strategy driven by a recovery-type error estimator[34], which significantly reduce the cost and improve the accuracy through dynamically focusing the mesh resolution on the fracture region. The main solver includes a robust Newton-Raphson scheme for the strongly coupled displacement-phase-field system. Inherited from FEALPy, GPU acceleration is ready for large-scale 3D simulations, and AD is also available to skip conventional matrix assembly in rapid tests for new models.
@@
6 Discussion and Conclusions
This work introduced FEALPy, a tensor-centric numerical simulation engine that unifies traditional numerical simulation methods with modern tensor computation/DL frameworks. The open-source nature of FEALPy, along with its modular design, allows for easier contributions from the community.
By organizing computations around a consistent tensor abstraction layer, FEALPy achieves interoperability across multiple backends such as NumPy, PyTorch, and JAX, allowing seamless execution on diverse hardware architectures including CPUs and GPUs. This design not only allows us researchers to leverage the latest advances in hardware acceleration for tensors but also supports a wide range of software ecosystems.
The tensor abstraction layer also results in the backend independence, which allows developers to integrate future improvements. Users and developers can make FEALPy adopt to new tensor libraries or user C/C++ extensions for specific hardware platforms, by simply adding a new backend adapter. Newly-added tensor backends are available for all existing build-in/user algorithms; while new algorithms can run on all existing backends without modification (Unless the algorithm takes advantage of a tensor library’s unique features).
The modular architecture of FEALPy—comprising the tensor, commonality, algorithmic, and field layers—effectively separates mathematical abstraction from computational realization. This design provides both clarity and flexibility for researchers to implement, extend, and combine numerical algorithms without modifying the underlying data structures.
@@ 6 Discussion and Conclusions
All of these features together lower the migration barrier and learning cost for researchers to adopt and contribute to FEALPy, fostering a collaborative environment for advancing numerical simulation methods and their applications. We have been expanding the collection of educational examples and organizing the documentation of FEALPy to make it more accessible for students and researchers new to the field, encouraging contributions from the community.
@@
Source code of numerical examples are available at https://github.com/weihuayi/fealpy/tree/master/example. The user manual for FEALPy: http://suanhai.com.cn/intro.html.